# Beyond NISQ: The Megaquop Machine

### v1

@@ Abstract.
This article is a lightly edited transcript of my keynote address at the Q2B 2024 Conference in Silicon Valley on 11 December 2024.
Submitted to ACM Transactions on Quantum Computing.
quantum computation, quantum error correction and fault tolerance
ccs: Hardware Quantum computationccs: Hardware Quantum error correction and fault toleranceccs: Hardware Quantum computationccs: Hardware Quantum error correction and fault tolerance
@@ 2. The road to fault tolerance
A snapshot of the state of the art comes from the Google Quantum AI team (Acharya et al., 2024). Their recently introduced Willow superconducting processor has improved transmon lifetimes, measurement errors, and leakage correction compared to its predecessor Sycamore (AI, 2023). With it they can perform millions of rounds of surface-code error syndrome measurement with good stability, each round lasting about a microsecond. Most notably, they find that the logical error rate per measurement round improves by a factor of 2 (a factor they call $\Lambda$) when the code distance increases from 3 to 5 and again from 5 to 7, indicating that further improvements should be achievable by scaling the device further. They performed accurate real-time decoding for the distance 3 and 5 codes. To further explore the performance of the device they also studied the repetition code, which corrects only bit flips, out to a much larger code distance. As the hardware continues to advance we hope to see larger values of $\Lambda$ for the surface code, larger codes achieving much lower error rates, and eventually not just quantum memory but also logical two-qubit gates with much-improved fidelity compared to the fidelity of physical gates.
@@ 4. Trading simplicity for performance
One approach uses cat qubits, in which the encoded 0 and 1 are coherent states of a microwave resonator, well separated in phase space, such that the noise afflicting the qubit is highly biased. Bit flips are exponentially suppressed as the mean photon number of the resonator increases, while the error rate for phase flips induced by loss from the resonator increases only linearly with the photon number. This year the AWS team built a repetition code to correct phase errors for cat qubits that are passively protected against bit flips, and showed that increasing the distance of the repetition code from 3 to 5 slightly improves the logical error rate (Putterman et al., 2024). (See also Ref. (Réglade et al., 2024).)
@@
Whether this trading of simplicity for performance in superconducting qubits will ultimately be advantageous for scaling to large systems is still unclear. But it’s appropriate to explore such alternatives which might pay off in the long run.
@@ 6. Toward the megaquop machine
How can we reach the megaquop regime? More efficient quantum codes like those recently discovered by the IBM team might help (Bravyi et al., 2024). These require geometrically nonlocal connectivity and are therefore better suited for Rydberg optical tweezer arrays than superconducting processors, at least for now. Error mitigation strategies tailored for logical circuits might help by boosting the circuit volume that can be simulated beyond what one would naively expect based on the logical error rate. Recent advances from the Google team, which reduce the overhead cost of logical gates, might also be helpful (Gidney et al., 2024).
@@
As a scientist, I’m intrigued by what we might conceivably learn about quantum dynamics far from equilibrium by doing simulations on megaquop machines, particularly in two spatial dimensions. But when seeking quantum advantage in that arena we should bear in mind that classical methods for such simulations are also advancing impressively. (See, for example, Ref. (Begušić and Chan, 2024) and (Angrisani et al., 2024).)
@@ Acknowledgements.
This article is based on a keynote address delivered at the Q2B 2024 Conference in Silicon Valley. I thank Matt Johnson for inviting me to speak at Q2B for each of the past eight years, and Travis Humble for encouraging me to publish the talk. I also thank Dorit Aharonov, Sergio Boixo, Earl Campbell, Roland Farrell, Ashley Montanaro, Mike Newman, Will Oliver, Chris Pattison, Rob Schoelkopf, and Qian Xu for helpful comments. I gratefully acknowledge support from the US Department of Energy and the National Science Foundation. The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.

### v2

@@ Abstract.
Keynote address at the Q2B 2024 Conference in Silicon Valley on 11 December 2024.
quantum computing, quantum error correction, fault-tolerant quantum computing.
doi: 10.1145/3723153journal: TQCjournalyear: 2025publicationmonth: 3ccs: Hardware Quantum computationccs: Hardware Quantum error correction and fault tolerance
@@ 2. The road to fault tolerance
A snapshot of the state of the art comes from the Google Quantum AI team (Acharya et al., 2025). Their recently introduced Willow superconducting processor has improved transmon lifetimes, measurement errors, and leakage correction compared to its predecessor Sycamore (Acharya et al., 2023). With it they can perform millions of rounds of surface-code error syndrome measurement with good stability, each round lasting about a microsecond. Most notably, they find that the logical error rate per measurement round improves by a factor of 2 (a factor they call $\Lambda$) when the code distance increases from 3 to 5 and again from 5 to 7, indicating that further improvements should be achievable by scaling the device further. They performed accurate real-time decoding for the distance 3 and 5 codes. To further explore the performance of the device they also studied the repetition code, which corrects only bit flips, out to a much larger code distance. As the hardware continues to advance we hope to see larger values of $\Lambda$ for the surface code, larger codes achieving much lower error rates, and eventually not just quantum memory but also logical two-qubit gates with much-improved fidelity compared to the fidelity of physical gates.
@@ 4. Trading simplicity for performance
One approach uses cat qubits, in which the encoded 0 and 1 are coherent states of a microwave resonator, well separated in phase space, such that the noise afflicting the qubit is highly biased. Bit flips are exponentially suppressed as the mean photon number of the resonator increases, while the error rate for phase flips induced by loss from the resonator increases only linearly with the photon number. The AWS team, using their recently announced Ocelot quantum chip, built a repetition code to correct phase errors for cat qubits that are passively protected against bit flips, and showed that increasing the distance of the repetition code from 3 to 5 slightly improves the logical error rate (Putterman et al., 2025). (See also Ref. (Réglade et al., 2024).)
@@
A theoretically compelling idea is to fashion an intrinsically robust qubit based on Majorana zero modes in superconducting nanowires. This scheme, tenaciously pursued by Microsoft, requires sophisticated fabrication and has progressed slowly so far. Their recently announced Majorana 1 processor incorporates fermion-parity readout circuitry which might pave the way for the first convincing demonstration of a topologically protected qubit (Aghaee et al., 2025).
Whether such trading of simplicity for performance in quantum devices will ultimately be advantageous for scaling to large systems is still unclear. But it’s appropriate to explore such alternatives which might pay off in the long run.
@@ 6. Toward the megaquop machine
How can we reach the megaquop regime? More efficient quantum codes like those recently discovered by the IBM team might help (Bravyi et al., 2024). These require geometrically nonlocal connectivity and are therefore better suited for Rydberg optical tweezer arrays than superconducting processors, at least for now. Error mitigation strategies tailored for logical circuits might help by boosting the circuit volume that can be simulated beyond what one would naively expect based on the logical error rate (Aharonov et al., 2024). Recent advances from the Google team, which reduce the overhead cost of logical gates, might also be helpful (Gidney et al., 2024).
@@
As a scientist, I’m intrigued by what we might conceivably learn about quantum dynamics far from equilibrium by doing simulations on megaquop machines, particularly in two spatial dimensions. But when seeking quantum advantage in that arena we should bear in mind that classical methods for such simulations are also advancing impressively. (See, for example, Ref. (begušić2024realtimeoperatorevolutiondimensions) and (Angrisani et al., 2024).)
@@ Acknowledgements.
This article is based on a keynote address delivered at the Q2B 2024 Conference in Silicon Valley. I thank Matt Johnson for inviting me to speak at Q2B for each of the past eight years, and Travis Humble for encouraging me to publish the talk. I also thank Dorit Aharonov, Jason Alicea, Sergio Boixo, Earl Campbell, Roland Farrell, Ashley Montanaro, Mike Newman, Will Oliver, Chris Pattison, Rob Schoelkopf, and Qian Xu for helpful comments. I gratefully acknowledge support from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator, and the National Science Foundation (PHY-2317110). The Institute for Quantum Information and Matter is an NSF Physics Frontiers Center.